OpenStudy (anonymous):

Solve the system of equations using substitution. 3y = -1/2x + 2 y = -x + 9

OpenStudy (help!!!!):

I would rearranged theses to x+y=numbers

OpenStudy (luigi0210):

Sub in y from the second equation into the first

OpenStudy (xapproachesinfinity):

you have y=-x+9 then 3y=-1/2x+2 becomes 3(-x+9)=-1/2x+2 -3x+27=-1/2x+2 solve for x